A car initially traveling at a speed of 16 meters per second accelerates uniformly to a speed of

20 meters per second over a distance of 36 meters.  What is the magnitude of the car’s acceleration?


1
Expert's answer
2022-03-27T10:55:47-0400

The magnitude of the car’s acceleration

"a=\frac{v_f^2-v_i^2}{2d}"

"a=\frac{20^2-16^2}{2*36}=2\:\rm m/s^2"